# Optimized Fish Locomotion using Design-by-Morphing and Bayesian Optimization

### v1

@@ 5 Conclusions
This work introduced a morphing-based parametric design framework, integrated with Bayesian optimization, to identify energetically efficient undulatory swimming profiles. By expressing the body deformation as a linear combination of diverse baseline shapes, the proposed method was able to explore a broad design space that includes both conventional and unconventional swimming kinematics. The optimization process reveals a best-performing profile with an efficiency of 82.4%, followed by two other high-performing profiles with efficiencies of 51.8% and 42.8%, each surpassing the performance of traditional anguilliform and carangiform profiles. Our results also explain that the optimal and second-best swimming profiles are closely aligned with the anguilliform swimming mode, with the distinction that the anterior motion exhibits a phase lag. It can be concluded that the nature-inspired anguilliform swimming mode, with modifications introduced through the proposed optimization strategy, leads to superior performance.
@@ Acknowledgements
MSU Khalid acknowledges the funding support from the Natural Sciences and Engineering Research Council of Canada (NSERC) through the Discovery grant program. The computational work and simulations reported here were performed on the supercomputing clusters administered and managed by the Digital Research Alliance of Canada.

### v2

@@ 5 Conclusions
This work introduced a morphing-based parametric design framework, integrated with Bayesian optimization, to identify energetically efficient undulatory swimming profiles. By expressing the body deformation as a linear combination of diverse baseline shapes, the proposed method was able to explore a broad design space that includes both conventional and unconventional swimming kinematics. The optimization process reveals a best-performing profile that attains a peak propulsive efficiency of approximately $57\%$ at a low undulation frequency of $f=0.2$ and maintains high efficiency levels between $49\%$ and $57\%$ over a broad region of the kinematic space, corresponding to an overall enhancement of roughly $16\%$$35\%$ compared to conventional anguilliform and carangiform swimming kinematics. Our results also explain that the optimal profile is closely aligned with the anguilliform swimming mode, with the distinction that the anterior motion exhibits a phase lag. It can be concluded that the nature-inspired anguilliform swimming mode, with modifications introduced through the proposed optimization strategy, leads to superior performance.
@@ Acknowledgements
MSU Khalid acknowledges the funding support from the Natural Sciences and Engineering Research Council of Canada (NSERC) through the Discovery grant program. The computational work and simulations reported here were performed on the supercomputing clusters administered and managed by the Digital Research Alliance of Canada. The authors would also like to acknowledge the invaluable comments regarding arc-length preservation from a peer-reviewer on an earlier draft of this paper, which led to significant improvement of our work.